OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4 PLEASE HELP I KEEP GETTING IT WRONG

OpenStudy (anonymous):

no you have the -4 also

OpenStudy (anonymous):

its x+3 = x-1 +2

OpenStudy (anonymous):

you end up with no solution even if you distribute -2x...=-2x...

mathslover (mathslover):

\[\large{-2(x+3)=-2(x+1)}\] \[\large{\frac{-2(x+3)}{-2}=\frac{-2(x+1)}{-2}}\] \[\large{\frac{\cancel{-2}^1(x+3)}{\cancel{-2}^1}=\frac{\cancel{-2}^1(x+1)}{\cancel{-2}^1}}\] \[\large{x+3=x+1}\] \[\large{x+3-x=1}\] \[\large{3\ne 1}\] Hence no solution

hero (hero):

@mathslover, you were just telling me that I forgot the -4 and you did it too.

OpenStudy (anonymous):

or infinite -2x-6 = -2x-2-4 = -2x-6

OpenStudy (theviper):

@mathslover -4 is included

OpenStudy (theviper):

\[\Huge{\color{red}{\implies -2x-6=-2x-2-4}}\]\[\LARGE{\color{blue}{\implies \cancel{-2x}-6=\cancel{-2x}-2-4}}\]\[\LARGE{\color{green}{\implies-6=-6}}\]It's a true statement :)

OpenStudy (theviper):

@TuringTest help plz :)

hero (hero):

Not necessary. You got it correct Viper

OpenStudy (theviper):

ummm ok thanx brother @Hero :)

hero (hero):

of course, he gave you the medal for the correct answer. I messed up because I didn't see the -4

OpenStudy (theviper):

but why I m there U deserve a medal ;)

hero (hero):

If I'd seen the -4, I would have said to divide both sides by -2 to get x + 3 = x + 1 + 2 x + 3 = x + 3

mathslover (mathslover):

Oops sorry for that mistake @Hero and @TheViper

OpenStudy (theviper):

np ur first mistake ;)